OpenStudy (anonymous):

Having a lot of trouble with this for some reason... Find the 9th term of the arithmetic sequence. a7= -42 and a11= -28

OpenStudy (anonymous):

Its that flvs precal?

Parth (parthkohli):

Looks like 3.5 is being added to get the subsequent terms

Parth (parthkohli):

a7 = -42 a8 = -42 + 3.5 a9 = -42 + 3.5 + 3.5

OpenStudy (jdoe0001):

ditto

OpenStudy (anonymous):

Ah! Thank you so much @ParthKohli! So a9 would equal -35

Parth (parthkohli):

Great, you can also use many formulas for problems like these. I have an alternative method to do this.

OpenStudy (jdoe0001):

you'd just have to bear in mind that to get the common ratio from the 7th term to the 11th term, you'd need to use it 4 times so \(\bf \textit{common ratio}=\cfrac{-28-(-42)}{4}\)

Parth (parthkohli):

If you are given two terms, and want to calculate the middle term, take the average of the two terms.

Parth (parthkohli):

In this problem, a9 happens to fall right in the middle of a7 and a11.

OpenStudy (anonymous):

Thank you so much! The formula we were given to solve this an= a1+(n-1)d just was not making much sense to me.

Parth (parthkohli):

You can use that one too! Actually, that is what we generally use for problems like these.

Parth (parthkohli):

Do ask if you need the general way.